Construct the indicated confidence interval for the population mean

muμ using the​ t-distribution. Assume the population is normally distributed.

cequals=0.99 x overbarxequals=13.6​, sequals=0.56 nequals=15

Round to one decimal place as needed
Answer #1

Given that,

\bar x = 13.6

s =0.56

n = 15

Degrees of freedom = df = n - 1 = 15- 1 = 14

At 99% confidence level the t is ,

\alpha = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

t\alpha /2  df = t0.005,14 = 2.977

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.977* (0.56 / \sqrt 15) = 0.4

The 99% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

13.6 - 0.4 < \mu < 13.6+ 0.4

13.2 < \mu < 14.0

(13.2 ,14.0 )